# BOTH2Hands: Inferring 3D Hands from Both Text Prompts and Body Dynamics

### 20 Dec 2023 v3

@@ 4.3 Cross-attention hand blending
Inspired by the success of the sharing-weights transformer in InterGen [25], we adopted a cross-attention transformer for the gestures blending part. The networks are fed with two conditioned hand motions, $\textbf{X}_{T}$ and $\textbf{X}_{B}$, we sequentially apply these as attention inputs to the transformer, and compute the weighted reconstruction loss between the final output and two types of gestures. Specifically, the $\textbf{X}_{T}$ and $\textbf{X}_{B}$ are firstly embedded into a common latent space and positionally encoded into the latent states $\textbf{h}_{text}$ and $\textbf{h}_{body}$. Then, it is processed by $N$ attention-based blocks to obtain the blending hidden states $\textbf{h}^{N}_{out}$. Each block consists of multi-head cross-attention layers ($Attn$) followed by one feed-forward network ($FF$). For first-time hands passing through cross-attention block, the input hidden layer $\textbf{h}_{body}$ is embedded into the query matrix (Q), the attention hidden layer $\textbf{h}_{text}$ are embedded into a key matrix (K) and value matrix (V), finally embed results into a vector $\textbf{h}^{(1)}_{out}$, the computation process is detailed below:
@@
where $D$ is the number of channels in the attention layer, $W$ are trainable weights, and $\textbf{Q}^{(i)}$, $\textbf{K}^{(i)}$, $\textbf{V}^{(i)}$ are transformer matrixes under $i-th$ layer. Passing through the attention layer once, we get the attention layer out $\textbf{h}^{(1)}_{out}$, then we switch the K, V input to $\textbf{h}_{body}$, which means we use body conditioned hands as attention input to emphasizing the body physical movements. The changed attention process is:
@@
We use blending loss to supervise weights $W$ learning.
$\mathcal{L}_{blend}=\mathbb{E}_{{X}_{S},{X}_{B},{X}_{T}}||X_{S}-(w_{B}X_{B}+w_{T}X_{T})||_{1},$ (13)
where $w_{B}$ and $w_{T}$ are hyperparameters controlling weights of different hand motion parts, we set them all positive numbers and $w_{B}+w_{T}=1$.
@@ 5 Experiment
Aiming to evaluate the validity of our method and dataset, we design various experiments. For method evaluation, we compare our approach and baseline with existing methods in human motion synthesis (Sec. 5.1). To validate the richness and effectiveness of the BOTH57M, we train our method separately on training sets of BOTH57M and Motion-X and subsequently evaluated through their test sets (Sec. 5.2). An ablation study is also performed to verify the importance of hand projection and blending loss (Sec. 5.3).
@@
T2M-GPT ${\color[rgb]{1,0,0}0.042}^{\pm.014}$ $0.073^{\pm.014}$ $0.104^{\pm.020}$ $0.461^{\pm.016}$ $1.398^{\pm.010}$ $3.689^{\pm.094}$ $1.178^{\pm.100}$
MDM $0.039^{\pm.020}$ ${\color[rgb]{1,0,0}0.077}^{\pm.016}$ ${\color[rgb]{0,0,1}0.114}^{\pm.024}$ $0.257^{\pm.024}$ $1.397^{\pm.008}$ $3.887^{\pm.074}$ $1.273^{\pm.086}$
@@
BOTH2Hands (Ours) $0.037^{\pm.014}$ $0.075^{\pm.020}$ ${\color[rgb]{1,0,0}0.115}^{\pm.028}$ ${\color[rgb]{0,0,1}0.201}^{\pm.020}$ ${\color[rgb]{0,0,1}1.392}^{\pm.008}$ $3.969^{\pm.082}$ ${\color[rgb]{1,0,0}1.312}^{\pm.034}$
BOTH2Hands-Text $0.035^{\pm.020}$ $0.067^{\pm.026}$ $0.109^{\pm.030}$ ${\color[rgb]{1,0,0}0.198}^{\pm.012}$ ${\color[rgb]{1,0,0}1.391}^{\pm.006}$ $3.980^{\pm.090}$ ${\color[rgb]{0,0,1}1.274}^{\pm.138}$
BOTH2Hands-Body ${\color[rgb]{0,0,1}0.039}^{+.012}$ ${\color[rgb]{0,0,1}0.076}^{\pm.024}$ $0.112^{\pm.026}$ $0.203^{\pm.016}$ $1.392^{\pm.010}$ $3.955^{\pm.098}$ $1.266^{\pm.120}$
@@ 5.1 Methods Evaluation
We benchmark our approach against various other methods on the task of using text and body conditions to generate hands. We introduced latent text-to-motion methods T2M-GPT [67] and MLD [6], diffusion-based method MDM [56], and body-conditioned motion synthesis method EgoEgo [22]. For unbiased comparison, we aligned the input and output dimensions, keeping text conditions unchanged and custom-designed for body conditions. In MLD, we employ two encoder-decoder structures for the body and hands. We merge the cleaned latent body token with the noisy hand token in latent diffusion, denoise the combined token for latent predicted hands, and feed it into the hand decoder. For T2M-GPT, we trained an encoder-decoder structure to derive body features and then add up body and text tokens. For non-latent space methods, we directly concatenate the body conditions onto the noised hands and treat text conditions as MDM [56]. All other methods use the same implementation details as they presented, for structures we add, keeping dimensions the same as the original framework. For our method, all transformers consist of N=4 blocks, a latent dimension of 512, and 4 attention heads. We use a frozen CLIP-ViTL-14 model as the text encoder. During training, the diffusion timesteps are set to 1000 during training and inference, using the AdamW optimizer with a fixed learning rate of $1e^{-4}$, and hyperparameters $w_{B}=0.8,w_{T}=0.2$, the motion blending is performed 3 times. All the methods are trained with the BOTH57M training set on a single NVIDIA GeForce RTX 2080 Ti GPU for about 2 Days. For more inference results, please refer to supplementary materials.
@@
Motion-X Motion-X $0.048^{\pm.018}$ $0.364^{\pm.040}$ $3.203^{\pm.116}$ $1.087^{\pm.072}$
@@
BOTH57M $0.037^{\pm.014}$ $0.201^{\pm.020}$ $3.969^{\pm.082}$ $1.312^{\pm.034}$
Table 3: Cross-dataset comparisons of BOTH57M and Motion-X. We train our pipeline on the training set of them, then evaluate the models on their test sets. We use a 95% confidence interval, approximated by the mean value plus or minus twice the standard deviation.
The evaluation metrics are composed of the following sections: Motion-retrieval precision (R Precision). Fréchet Inception Distance (FID) [17]. Multi-modal Distance (MM-Dist). Diversity and MultiModality (MModality). Tab. 2 presents detailed quantitative results from the same test set, showing our methods reconstructed motions closest to the real motion. Fig. 4 shows BOTH2Hands achieves good alignment between hand motion and conditions. Non-latent methods perform well on body, but poorly on text conditions. Latent methods struggle with body conditions. Fig. 6 demonstrates our blending block beats our diffusion baseline. Text-conditioned hands lack body conditions, while body-conditioned hands fail to meet prompt requirements.
@@ 5.2 Dataset Evaluation
To showcase the richness of the hand motions and the accuracy of text-guided actions, we conduct a comparison between our BOTH57M and Motion-X [26], the largest full-body motion dataset currently featuring text annotations. We train our BOTH2Hands method on training sets Motion-X and BOTH57M separately. Then perform validation on respective test sets. The comparison results are presented in Tab. 3, BOTH57M provides better alignment from text to hands, while hand diversity is also better than Motion-X. Fig. 5 shows our qualitative results, given body and text condition on test sets, BOTH57M trained method always perform better alignment on text and body condition, for general motion prompts also performs well in case we have general motion annotates that contains hand description.
Figure 6: Qualitative results of baseline comparison. The motions follow texts are circled red, the motion follow body is circled blue.
Figure 7: Qualitative results of ablation study. The motions errors are circled red.
@@ 5.3 Ablation Study
To validate the importance of hand projection and blending loss, we conduct an ablation study, assessing the impact of removing these components from the model. As loss replace, we randomly choose fingers on hands as the target, using linear distance loss. Numerical results in Tab. 4 indicate our method enhanced performance with the inclusion of hand projection, allowing the transformer to focus solely on the motion, thereby negating the impact of the reference system. Blending loss highly improves hand motion quality, proving that learning the hand as a whole is useful for keeping features of hand motions. Fig. 7 presents our inference outcomes. Without hand projection and blending loss, we observe prominent finger errors.
@@ 6 Conclusion
We introduced BOTH57M, the first comprehensive body-hand dataset that incorporates precise gestures and body movements, paired with meticulous finger-level hand annotations and body descriptions, which spans a variety of activities, consisting of 57.4 million frames in 8.31 hours, supplemented with 23,477 text annotations. Based on this dataset, we presented BOTH2Hands, a robust algorithm initiating procedure for the generation of hand movements under two conditions, such as body movements and text prompts, and then utilized a cross-attention transformer for motion blending. We also conducted a series of detailed evaluations to demonstrate the robustness of our methods and the enhancement of our dataset for the two-hand generation task. We believe the BOTH57M could boost future exploration in multi-modal control and the analysis of human behaviour.
@@ A More Details of BOTH57M
In this section, we delve deeper into the specifics of our data collection system and the process of annotation.
@@ A.1 Camera Dome Setting
Our RGB cameras are mostly arranged evenly in three layers(low, medium, and high) at 270°. The rest of the cameras were arranged on the highest layer at the remaining 90°. We zoom in on some of the cameras in order to ensure that they are sufficiently fine so that the motion of hands can be captured in detail. Moreover, we placed three 5500w fill lights in the dome to support hand motion capture with high-precision zoom which highly requires brightness. Fig. 8 shows the cameras and fill lights setting in the dome, the red cameras are zoomed in for capturing hands.
@@ A.2 Capture System Synchronization
In order to avoid motion blurring, the RGB cameras are set to work at 59.97 FPS with the resolution adjusted to 3840$\times$2160. Furthermore, all cameras will not get ready until all the cameras are turned on and set to the intended focal length. When capturing, everyone is required to make a T-pose at the beginning and end of each shoot so that the initial orientation alignment of ground truth can be easily performed, i.e. the origin of the coordinate system is at the initial position of the person and the forward direction is towards the z-coordinate with up direction is towards the y-coordinate, using a right-handed system. Fig. 9 provides the gallery of data examples captured by our dome with 32 synchronized high-resolution RGB cameras.
A.3 Data Process
@@
After reconstructing the human skeleton, we offer the data together with the book ”Dictionary of Gestures” [5] to different annotators. For each motion sequence, three annotators are supposed to describe the motion given according to the emotion from the corresponding poses in the book in ”Physical Description + Meaning of the Action” format, where ”Physical Description” means to describe the action precisely and ”Meaning of the Action” means the meaning the action wants to express. For example, ”Put the five fingers of your right hand together in front of your right ear” is a physical description, and ”Such an action is meant to express listening carefully” is a meaning of the action. The other three annotators are required to focus on only hand movements and use detailed descriptions to explain how the figures move. The annotations contain only physical descriptions such as ”Left-hand index finger straight, other fingers are naturally bent, right-hand thumb extended, other fingers in a fist”. Fig. 10 shows the motion of our dataset as well as the corresponding annotation in detail.

### 09 Apr 2024 v4

@@ 4.3 Cross-attention hand blending
Inspired by the success of the sharing-weights transformer in InterGen [29], we adopted a cross-attention transformer for gesture blending. The networks are fed with two conditioned hand motions, $\textbf{X}_{T}$ and $\textbf{X}_{B}$. We sequentially apply hand motions as attention inputs to the transformer, and compute the weighted reconstruction loss between the final output and two types of gestures. Specifically, the $\textbf{X}_{T}$ and $\textbf{X}_{B}$ are firstly embedded into a common latent space and positionally encoded into the latent states $\textbf{h}_{text}$ and $\textbf{h}_{body}$. Then, it is processed by $N$ attention-based blocks to obtain the blending hidden states $\textbf{h}^{N}_{out}$. Each block consists of multi-head cross-attention layers ($Attn$) followed by one feed-forward network ($FF$). For the first time the hands passing through the cross-attention block, the input hidden layer $\textbf{h}_{body}$ is embedded into the query matrix (Q); the attention hidden layer $\textbf{h}_{text}$ is embedded into a key matrix (K) and value matrix (V); finally we embed results into a vector $\textbf{h}^{(1)}_{out}$. The hand-blending process is detailed below:
@@
where $D$ is the number of channels in the attention layer; $W$ are trainable weights, and $\textbf{Q}^{(i)}$, $\textbf{K}^{(i)}$, $\textbf{V}^{(i)}$ are transformer matrices under i-th layer. Passing through the attention layer once, we get the output $\textbf{h}^{(1)}_{out}$. Then we switch the K, V input to $\textbf{h}_{body}$, which means we use body-conditioned hands as attention input to emphasize the body movements. The changed attention process is:
@@
We use blending loss to supervise the learning of weights $W$.
$\mathcal{L}_{blend}=\mathbb{E}_{{X}_{GT},{X}_{B},{X}_{T}}||X_{GT}-(w_{B}X_{B}+w_{T}X_{T})||_{1},$ (13)
where $w_{B}$ and $w_{T}$ are hyperparameters controlling weights of different hand motion parts. We set $w_{B}$ and $w_{T}$ to positive numbers and $w_{B}+w_{T}=1$. $X_{GT}$ is GT hand motion.
@@ 5 Experiment
We design various experiments to evaluate the validity of our method and dataset. For method evaluation, we compare our approach and baseline with existing human motion synthesis methods (Sec. 5.1). To assess the richness and effectiveness of the BOTH57M, we train our method on training sets of BOTH57M and Motion-X separately. And subsequently evaluated trained models through the test sets (Sec. 5.2). Additionally, an ablation study is performed to verify the importance of hand projection and blending loss (Sec. 5.3).
@@
T2M-GPT ${\color[rgb]{1,0,0}\textbf{0.042}}^{\pm.014}$ $0.073^{\pm.014}$ $0.104^{\pm.020}$ $0.461^{\pm.016}$ $1.398^{\pm.010}$ $3.689^{\pm.094}$ $1.178^{\pm.100}$
MDM $0.039^{\pm.020}$ ${\color[rgb]{1,0,0}\textbf{0.077}}^{\pm.016}$ ${\color[rgb]{0,0,1}\textbf{0.114}}^{\pm.024}$ $0.257^{\pm.024}$ $1.397^{\pm.008}$ $3.887^{\pm.074}$ $1.273^{\pm.086}$
@@
BOTH2Hands (Ours) $0.037^{\pm.014}$ $0.075^{\pm.020}$ ${\color[rgb]{1,0,0}\textbf{0.115}}^{\pm.028}$ ${\color[rgb]{0,0,1}\textbf{0.201}}^{\pm.020}$ ${\color[rgb]{0,0,1}\textbf{1.392}}^{\pm.008}$ $3.969^{\pm.082}$ ${\color[rgb]{1,0,0}\textbf{1.312}}^{\pm.034}$
BOTH2Hands-Text $0.035^{\pm.020}$ $0.067^{\pm.026}$ $0.109^{\pm.030}$ ${\color[rgb]{1,0,0}\textbf{0.198}}^{\pm.012}$ ${\color[rgb]{1,0,0}\textbf{1.391}}^{\pm.006}$ $3.980^{\pm.090}$ ${\color[rgb]{0,0,1}\textbf{1.274}}^{\pm.138}$
BOTH2Hands-Body ${\color[rgb]{0,0,1}\textbf{0.039}}^{+.012}$ ${\color[rgb]{0,0,1}\textbf{0.076}}^{\pm.024}$ $0.112^{\pm.026}$ $0.203^{\pm.016}$ $1.392^{\pm.010}$ $3.955^{\pm.098}$ $1.266^{\pm.120}$
@@ 5.1 Methods Evaluation
We compare our approach with several other methods in the task of generating hands based on textual and body conditions. We introduce latent text-to-motion methods T2M-GPT [74] and MLD [6], diffusion-based method MDM [62], and body-conditioned motion synthesis method EgoEgo [25] for comparison. We align the input and output dimensions for unbiased comparison, keeping text and body conditions the same for all methods. In MLD, we employ two encoder-decoder structures for the body and hands. In the latent diffusion process, we merge the cleaned latent body token with the noisy hand token. The combined token is then denoised to predict latent hands, which are subsequently fed into the hand decoder. For T2M-GPT, we train an encoder-decoder structure to derive body features and then add up body and text tokens. For non-latent space methods, we directly concatenate the body conditions onto the noised hands as input and follow [62] to add text conditions into it. All methods use the same implementation details as they presented. For the structures added to other methods, we keep dimensions the same as the original framework. For our method, all transformers consist of N=4 blocks, a latent dimension of 512, and 4 attention heads. We use a frozen CLIP-ViTL-14 model as the text encoder. As for other parameters, the diffusion timesteps are set to 1000 during training and inference; the AdamW optimizer is used with a fixed learning rate of $1e^{-4}$; and hyperparameter $w_{B}$ is set to 0.8, $w_{T}$ is set to 0.2; the motion blending process (method in Sec.4.3) is performed 3 times. All the methods are trained with the BOTH57M training set on a single NVIDIA GeForce RTX 2080 Ti GPU for about 2 days. For more inference results, please refer to supplementary materials.
@@
Motion-X Motion-X $\textbf{0.048}^{\pm.018}$ $\textbf{0.364}^{\pm.040}$ $3.203^{\pm.116}$ $\textbf{1.087}^{\pm.072}$
@@
BOTH57M $\textbf{0.037}^{\pm.014}$ $\textbf{0.201}^{\pm.020}$ $3.969^{\pm.082}$ $\textbf{1.312}^{\pm.034}$
Table 3: Cross-dataset comparisons of BOTH57M and Motion-X. We train our pipeline on the training set of them, then evaluate the models on their test sets. Real(GT) means the GT data in the training set is used for evaluation. We use a 95% confidence interval, approximated by the mean value plus or minus twice the standard deviation.
Following  [14], our evaluation metrics include Motion-retrieval precision (R Precision), Fréchet Inception Distance (FID) [19], Multi-modal Distance (MM-Dist), Diversity and MultiModality (MModality). And we randomly split BOTH57M into the train (80%), val (5%), and test (15%) set, adopting SMPLH as motion representation. Tab. 2 presents detailed quantitative results from the same test set, showing our methods reconstructed motions closest to the real motion. Fig. 4 shows BOTH2Hands achieves good alignment between hand motions and conditions. Non-latent methods perform well on body conditions, but poorly on text conditions. Latent methods struggle with body conditions. Fig. 6 demonstrates our blending block beats our diffusion baseline. Text-conditioned hands lack body alignment, while body-conditioned hands fail to meet prompt requirements. Nevertheless, we also see marginal improvements in our evaluation results. This phenomenon implies that multi-conditioned generation performance may not be adequately reflected by widely used metrics such as R-precision, which are effective for single-conditioned generation evaluation. Two key reasons are listed below. First, in a full-body setting, hand motion constitutes a small portion, limiting the metric-based improvements. Second, the correlation between hand motions and text is non-linear, the metrics increase brought by enhancing hand motion is limited due to complicated hand-text alignment. We plan to explore more suitable metrics for future studies and hope our released code and dataset can serve as a solid foundation for such exploration.
@@ 5.2 Dataset Evaluation
To highlight the richness of the hand motions and the accuracy of text prompts, we compare our BOTH57M with Motion-X [30], the largest full-body motion dataset with text currently. We train BOTH2Hands method on training sets of Motion-X and BOTH57M separately. Then validate methods on respective test sets. The comparison results are presented in Tab. 3. We add the GT data in the training set to the evaluation as the standard. The model trained on Motion-X training set performs well on the test set. However, the model trained on BOTH57M provides better alignment from text to hands, and its hand diversity is also better than the model trained on Motion-X. Fig. 5 shows our qualitative results. Given body and text conditions on test sets, our method trained on BOTH57M always performs better on text and body conditions. It also performs well on general motion prompts due to general motion annotations that contain hand descriptions in BOTH57M.
Figure 6: Qualitative results of baseline comparison. The motions following texts are circled red, the motions following body are circled blue.
Figure 7: Qualitative results of ablation study. The errors of motions are circled red.
@@ 5.3 Ablation Study
To validate the importance of hand projection and blending loss, we perform an ablation study by evaluating the effects of excluding these elements from the BOTH2Hands algorithm. Hand projection can be removed directly. But for blending loss, we choose linear distance loss as an alternative. Numerical results in Tab. 4 indicate our method performs better with hand projection. This process allows the transformer to focus solely on the motion. Blending loss also highly improves hand motion quality, proving that learning the hand from previous output is effective. As mentioned in Sec. 5.2, marginal improvements found in the ablation study also suffer from the insensitivity of the existing metrics used. But obvious improvements shown in Fig. 7 still prove the effectiveness of our hand projection and blending loss.
@@ 6 Conclusion
We introduce BOTH57M, the first comprehensive body-hand dataset that incorporates precise gestures and body movements, paired with meticulous finger-level hand annotations and body descriptions, which spans a variety of activities, consisting of 57.4 million frames in 8.31 hours, supplemented with 23,477 text annotations. Based on this dataset, we introduce BOTH2Hands, a robust algorithm designed to generate hand movements under two conditions: body movements and text prompts. Subsequently, we employ a cross-attention transformer for motion blending. We also conduct a series of detailed evaluations to demonstrate the robustness of our methods and the enhancement of our dataset for the two-hand generation task. We believe the BOTH57M could boost future exploration in multi-modal control and the analysis of human behavior.
7 Acknowledgement
This work was supported by National Key R&D Program of China (2022YFF0902301), Shanghai Local college capacity building program (22010502800). We also acknowledge support from Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI).
@@ A More Details of BOTH57M
In this section, we delve deeper into the specifics of our dataset. The following paragraphs introduce the camera dome setting, data reconstruction pipeline with data quality discussion, data text annotation, and data statics.
@@ A.1 Camera Dome Setting
Our RGB cameras are mostly arranged evenly in three layers(low, medium, and high) at 270°. Other cameras are arranged on the highest layer at the remaining 90°. We zoom in on some of the cameras in order to ensure that they are sufficiently fine so that the motion of hands can be captured in detail. Moreover, we placed three 5500W fill lights in the dome to support hand motion capture with high-precision zoom that highly requires brightness. Fig. 8 shows the cameras and fill lights setting in the dome, the red cameras are zoomed in for capturing hands.
@@ A.2 Capture System Synchronization
In order to avoid motion blurring, the RGB cameras are set to work at 59.97 FPS with the resolution adjusted to 3840$\times$2160. Furthermore, cameras are available only if they are all turned on and set to the intended focal length. When being captured, everyone is required to make a T-pose at the beginning and end of each shoot so that the initial orientation alignment of ground truth can be easily performed, i.e. the origin of the coordinate system is at the initial position of the person and the forward direction is towards the z-coordinate with up direction is towards the y-coordinate, using a right-handed system.
A.3 Actor Requirements
During shooting, actors should perform motions in “Dictionary of Gestures” [5] excluding unfriendly gestures. They are required to expand the pose in the book into a complete body-hand motion sequence and perform it. Performers are encouraged to increase the diversity of hand gestures and body motions without changing the original meaning of the book.
A.4 Data Reconstruction and Quality
For data reconstruction, we adopt ViTPose [66] for predicting 2D body keypoints, and MediaPipe [37] for estimating 2D hand keypoints from RGB images. Markerless mocap then derives SMPLH parameters and 3D coordinates from these 2D results. Fig. 9 provides an overlay of our dataset. We are confident that our dataset exhibits sufficient full-body quality in 32-camera views with partial camera zoom-in for hand motion. We quantitatively validated our dataset using a subset with manually labeled 2D keypoints, triangulated for 3D joint coordinates as a gold standard. Our annotations show 31.4 mm MPJPE and 25.2 mm PA-MPJPE overall, with hand metrics at 6.51 mm MPJPE and 3.97 mm PA-MPJPE, which is comparable to the widely adopted datasets, e.g., 36.1mm for Human3.6M [21] (full-body), 18.4mm for Grab [59] (body/hand marker-based), 2.78mm for InterHand [40] (hand only), 33.5mm for Motion-X [30] (pseudo-GT), etc. Nevertheless, we will release all the raw captured images, and welcome the community to further improve the tracking results.
Figure 9: Overlays of BOTH57M in SMPLH. Three different views are presented, arranged from left to right in order of temporal variation.
A.5 Data Text Annotation
For data text annotation, we segment each motion into clips of 5 seconds. We also offer the data together with the book “Dictionary of Gestures” [5] to the annotators. For each motion sequence, three annotators are supposed to describe the motion according to the emotion from the corresponding poses in the book in “Physical Description + Meaning of Action” format. “Physical Description” means to describe the action precisely; “Meaning of Action” means the emotion the action wants to express. For example, “Put the five fingers of your right hand together in front of your right ear.” is a physical description, and “Such an action is meant to express listening carefully.” is the meaning of the action. The other three annotators are required to focus on only hand movements and use detailed descriptions to explain how the fingers move. The annotations contain only physical descriptions such as “Left-hand index finger straight, other fingers are naturally bent, right-hand thumb extended, other fingers in a fist”. Fig. 10 shows the motion of our dataset as well as the corresponding annotation in detail.
@@ A.5 Data Text Annotation
Body part interact with # Frame Time
Head et al. 16.33M 2.33h
UpperBody et al. 5.64M 0.81h
Hand et al. 34.59M 5.01h
LowerBody et al. 1.27M 0.16h
Total 57.83M 8.31h
Table 5: BOTH57M Subjects.
A.6 Data Statics
BOTH57M is the only dataset that provides hybrid and detailed annotations of both body and hands at present. It consists of 1,384 motion clips and 57.4M frames, with 23,477 manually annotated motions and a rich vocabulary of 4,140 words. The data-capturing system includes 32 synchronized high-resolution RGB cameras with 59.97 FPS and 3840×2160 resolution, which are capable of full body capture tasks. As for subjects of the dataset, we follow “Dictionary of Gestures” to scientifically split the dataset into 36 subjects according to the body part that hands interact with, a statistic is shown in Tab.5.